# TerDiT: Ternary Diffusion Models with Transformers

### v1

@@ Abstract
Recent developments in large-scale pre-trained text-to-image diffusion models have significantly improved the generation of high-fidelity images, particularly with the emergence of diffusion models based on transformer architecture (DiTs). Among these diffusion models, diffusion transformers have demonstrated superior image generation capabilities, boosting lower FID scores and higher scalability. However, deploying large-scale DiT models can be expensive due to their extensive parameter numbers. Although existing research has explored efficient deployment techniques for diffusion models such as model quantization, there is still little work concerning DiT-based models. To tackle this research gap, in this paper, we propose TerDiT, a quantization-aware training (QAT) and efficient deployment scheme for ternary diffusion models with transformers. We focus on the ternarization of DiT networks and scale model sizes from 600M to 4.2B. Our work contributes to the exploration of efficient deployment strategies for large-scale DiT models, demonstrating the feasibility of training extremely low-bit diffusion transformer models from scratch while maintaining competitive image generation capacities compared to full-precision models. Code will be available at https://github.com/Lucky-Lance/TerDiT.
@@
Figure 1: Sample images (256$\times$256) generated by the ternary DiT model with 4.2B parameters. Comparison with full-precision diffusion transformer models (DiT-XL/2 with 675M parameters and Large-DiT-4.2B with 4.2B parameters) are shown in Fig. 4.
1 Introduction
@@
The advancements in large-scale pre-trained text-to-image diffusion models [1, 2, 3, 4, 5] have led to the successful generation of images characterized by both complexity and high fidelity to the input conditions. Notably, the emergence of diffusion models based on transformer architecture (DiTs) [6] represents a significant stride in this research domain. Compared with other diffusion models, diffusion transformers have demonstrated the capability to achieve lower FID scores with higher computation Gflops [6]. Recent research highlights the remarkable image generation capabilities of diffusion transformer architecture, as demonstrated in methods like Stable Diffusion 3 [7], and its impressive performance in video generation, showcased by work such as Sora111 https://openai.com/index/sora.
Given the impressive performance of diffusion transformer models, researchers are now increasingly delving into understanding the scaling law of these vision models [8], which resembles large language models (LLMs). For instance, the Stable Diffusion 3 provides trained DiT models across a spectrum of parameter sizes, from 800 million to 8 billion. Additionally, there is speculation among researchers that Sora might boast around 3 billion parameters. Given the enormous parameter numbers, deploying these DiT models can often be costly, especially on certain end devices.
@@
To deal with the deployment dilemma, there have already been some works on the efficient deployment of diffusion models recently [9, 10, 11, 12], most of which focus on model quantization. However, as far as we are concerned, there are still two main shortcomings in current research. Firstly, while much attention has been given to quantizing U-Net-based diffusion models, the exploration of quantization methods for transformer-based diffusion models remains quite limited. Secondly, most prevailing approaches in the current literature heavily rely on post-training quantization (PTQ) techniques for model quantization [9, 11, 13, 14, 15], which leads to unacceptable performance degradation, particularly with extremely low bit width (e.g., 2-bit and 1-bit). However, the extremely low-bit quantization of neural networks is important as it can significantly reduce the computation resources required for deployment [16], especially for models with huge parameter sizes. During our research, we find that there is still no work considering the extremely low-bit quantization of DiT models.
@@
To tackle these shortcomings, we propose to leverage the quantization-aware training (QAT) technique for the extremely low-bit quantization of large-scale DiT models. Low-bit QAT methods for large-scale models have been discussed in the LLM domain. Recent works observe that training large language models with extremely low-bit parameters (e.g., binary and ternary) from scratch can also lead to competitive performance [17, 18] compared with their full-precision counterparts. This indicates that there still exists significant precision redundancy in large-scale models, and implies the feasibility of the QAT scheme for large-scale DiT models.
@@
In this paper, we focus primarily on ternary weight networks [19] and provide TerDiT, the first quantization scheme for DiTs to our best knowledge. Our method achieves quantization-aware training (weight-only) and efficient deployment for ternary diffusion transformer models. Different from the naive quantization of linear layers in LLMs and CNNs [18, 19], we find that the direct weight ternarization of the adaLN module [20] in DiT blocks [6, 21] leads to large dimension-wise scale and shift values in the normalization layer compared with full-precision models (due to weight quantization, gradient approximation), which result in slower convergence speed and poor model performance. Consequently, we propose a variant of adaLN by applying an RMS Norm [22] after the ternary linear layers of the adaLN module to effectively mitigate this training issue.
@@
With this modification, we scale the parameters of the ternary DiT model from 600M (size of DiT-XL/2 [6]) to 4.2B (size of Large-DiT-4.2B [23]), finding that models with a larger number of parameters can converge to better results. We further employ existing 2-bit CUDA kernels to deploy the trained ternary DiT models, resulting in over a tenfold reduction in the model checkpoint size and about sixfold reduction in the inference memory consumption, while achieving competitive generation quality. The contributions are summarized as follows:
  1. 1.
@@
    Inspired by the quantization-aware training scheme for low-bit LLMs, we study the QAT methods for ternary DiT models and introduce DiT-specific improvements for better training, which have not been explored in DiT literature.
  2. 2.
@@
    We scale ternary DiT models from 600M to 4.2B parameters, and further deploy the trained ternary DiT on GPUs based on existing 2-bit CUDA kernels, enabling the inference of a 4.2B DiT model with less than 3GB GPU memory.
  3. 3.
@@
    Competitive evaluation results compared with full-precision models on the ImageNet [24] benchmark (image generation) showcase the effectiveness of our proposed TerDiT scheme.
@@
To our best knowledge, our study is the first attempt to explore the quantization of DiT models. We focus on quantization-aware training and efficient deployment for large ternary DiT models, offering valuable insights for future research on deploying DiT models with extremely low-bit precision.
2 Related Works
@@
Diffusion Models. Diffusion models have gained significant attention in recent years due to their ability to generate high-quality images and their potential for various applications. The concept of diffusion models was first introduced by [25], proposing a generative model that learns to reverse a diffusion process. This work laid the foundation for subsequent research in the field. [1] further extended the idea by introducing denoising diffusion probabilistic models (DDPMs), which have become a popular choice for image generation tasks. DDPMs have been applied to a wide range of domains, including unconditional image generation [1], image inpainting [26], and image super-resolution [27]. Additionally, diffusion models have been used for text-to-image synthesis, as demonstrated by the DALL-E model [28] and the Imagen model [5]. These models showcase the ability of diffusion models to generate highly realistic and diverse images from textual descriptions. Furthermore, diffusion models have been extended to other modalities, such as audio synthesis [29] and video generation [30], demonstrating their versatility and potential for multimodal applications.
Quantization of Diffusion Models. The quantization of diffusion models has been studied in recent years to improve the efficiency of diffusion models. Post-training quantization (PTQ) methods, such as those presented in [9, 11, 13, 14, 15], offer advantages in terms of quantization time and data usage. However, these methods often result in suboptimal performance when applied to low-bit settings. To address this issue, [31] proposes combining quantization-aware low-rank adapters (QALoRA) with PTQ methods, leading to improved evaluation results. As an alternative to PTQ, quantization-aware training (QAT) methods have been introduced specifically for low-bit diffusion model quantization [12, 32, 33, 34]. Despite their effectiveness, these QAT methods are currently only limited to small-sized U-Net-based diffusion models, revealing a research gap in applying QAT to large-scale DiT models. Further exploration of QAT techniques for large DiT models with extremely low bit width could potentially unlock even greater efficiency gains and enable the effective deployment of diffusion models in resource-constrained environments.
Ternary Weight Networks. Ternary weight networks [19] have emerged as a memory-efficient and computation-efficient network structure, offering the potential for significant reductions in inference memory usage. When supported by specialized hardware, ternary weight networks can also deliver substantial computational acceleration. Among quantization methods, ternary weight networks have garnered notable attention, with two primary approaches being explored: weight-only quantization and weight-activation quantization. In weight-only quantization, as discussed in [35], solely the weights are quantized to ternary values. On the other hand, weight-activation quantization, as presented in [36, 37], involves quantizing both the weights and activations to ternary values. Recent research has demonstrated the applicability of ternary weight networks to the training of large language models [18], achieving results comparable to their full-precision counterparts. Building upon these advancements, our work introduces, for the first time, quantization-aware training and efficient deployment schemes specifically designed for ternary DiT models. By leveraging the benefits of ternary quantization in the context of DiT models, we aim to push the boundaries of efficiency and enable the deployment of powerful diffusion models in resource-constrained environments, opening up new possibilities for practical applications.
3 TerDiT
In this section, we introduce TerDiT, a framework designed to conduct weight-only quantization-aware training and efficient deployment of large-scale ternary DiT models. We first give a brief review of diffusion transformer (DiT) models in Sec. 3.1. Then building upon the previous open-sourced Large-DiT [23], we illustrate the quantization function and quantization-aware training scheme in Sec. 3.2, conduct QAT-specific model structure improvement for better network training in Sec. 3.3, and introduce ternary deployment scheme in Sec. 3.4.
@@ 3 TerDiT
3.1 Diffusion Transformer Models
Diffusion Transformer. Diffusion transformer [6] (DiT) is an architecture that replaces the commonly used U-Net backbone in the diffusion models with a transformer that operates on latent patches. Similar to the Vision Transformer (ViT) architecture shown in Fig. 2 (C), DiT first pachifies the spatial inputs into a sequence of tokens, then the denoising process is carried out through a series of transformer blocks (Fig. 2 (B)). To deal with additional conditional information (e.g., noise timesteps $t$, class labels $l$, natural language inputs), DiT leverages adaptive normalization modules [38] (adaLN-Zero) to insert these extra conditional inputs to the transformer blocks. After the final transformer block, a standard linear decoder is applied to predict the final noise and covariance. The DiT models can be trained in the same way as U-Net-based diffusion models.
@@ 3.1 Diffusion Transformer Models
AdaLN is an important component in the DiT model [6] and has been proven more effective than cross-attention and in-context conditioning methods. Within the DiT architecture, the adaLN module integrates an MLP layer with a substantial number of parameters, constituting approximately 10% to 20% of the model’s total parameters. Throughout the training of TerDiT, we observe that the direct weight-ternarization of this module yields undesirable training results (analyzed in Sec. 3.3).
3.2 Model Quantization
As illustrated in Sec. 1, there is an increasing popularity in understanding the scaling law of DiT models, which has been proven crucial for developing and optimizing LLMs. In recent explorations, Large-DiT [23] successfully scales up the model parameters from 600M to 7B by incorporating the methodologies of LLaMA [39, 40] and DiT. The results demonstrate that parameter scaling can potentially enhance model performance and improve convergence speed for the label-conditioned ImageNet generation task. Motivated by this, we propose to further investigate the ternarization of DiT models, which can alleviate the challenges associated with deploying large-scale DiT models. In this subsection, we introduce the quantization function and quantization-aware training scheme.
@@ 3.2 Model Quantization
Referring to current popular quantization methods for LLMs [41, 42], we also multiply a learnable scaling parameter $\alpha$ to each ternary linear matrix after quantization, leading to the final value set as $\{-\alpha,0,+\alpha\}$. The quantization function is formulated as:
@@
$\text{RoundClip}(x,a,b)$ $=\text{Clamp}(\text{round}(x),a,b)\text{ and }\gamma=\frac{1}{mn}\sum_{ij}|W_{ij}|.$ (3)
TerDiT is a weight-only quantization scheme and we do not quantize the activations.
Quantization-aware Training Scheme. Based on the above-designed quantization function, we train a DiT model from scratch [^2] It is observed in [18] that for ternary LLMs, the conversion or post-training quantization from trained LLMs does not help. So we also train ternary DiT models from scratch. utilizing the straight-through estimator (STE) [43], allowing gradient propagation through the undifferentiable network components. We preserve the full-precision parameters of the network throughout the training process. For each training step, ternary weights are calculated from full-precision parameters by the ternary quantization function in the forward pass, and the gradients of ternary weights are directly applied to the full-precision parameters for parameter update in the backward pass.
@@
3.3 QAT-specific Model Structure Improvement
Activation Analysis for Ternary Linear Layer. In a ternary linear layer, all the parameters take one value from the set {-$\alpha$, 0, +$\alpha$}. The values passing through this layer would become large activation values, which might hamper the stable training of neural networks. We conduct a pilot study to qualitatively demonstrate the impact of ternary linear weights on activation values.

### v2

@@ Abstract
Recent developments in large-scale pre-trained text-to-image diffusion models have significantly improved the generation of high-fidelity images, particularly with the emergence of diffusion transformer models (DiTs). Among diffusion models, diffusion transformers have demonstrated superior image-generation capabilities, boosting lower FID scores and higher scalability. However, deploying large-scale DiT models can be expensive due to their excessive parameter numbers. Although existing research has explored efficient deployment techniques for diffusion models, such as model quantization, there is still little work concerning DiT-based models. To tackle this research gap, we propose TerDiT, the first quantization-aware training (QAT) and efficient deployment scheme for extremely low-bit diffusion transformer models. We focus on the ternarization of DiT networks, with model sizes ranging from 600M to 4.2B, and image resolution from 256$\times$256 to 512$\times$512. Our work contributes to the exploration of efficient deployment of large-scale DiT models, demonstrating the feasibility of training extremely low-bit DiT models from scratch while maintaining competitive image generation capacities compared to full-precision models. Our code and pre-trained TerDiT checkpoints have been released at https://github.com/Lucky-Lance/TerDiT.
Index Terms:
Model quantization, quantization-aware training, image generation, diffusion transformer models
@@
Figure 1: Sample images (256$\times$256) generated by TerDiT with 4.2B parameters (using 2GB of GPU memory) are shown. For comparison, images generated by full-precision diffusion transformer models—DiT-XL/2 with 675M parameters (using 3GB of GPU memory) and Large-DiT-4.2B with 4.2B parameters (using 17GB of GPU memory)—are provided in Fig. 5.
I Introduction
@@
The advancements in large-scale pre-trained text-to-image diffusion models [1, 2, 3, 4, 5, 6] have led to the successful generation of images characterized by both complexity and high fidelity to the input conditions. Notably, the emergence of diffusion transformer models (DiTs) [7] represents a significant stride in this research direction. Compared with other diffusion models, diffusion transformers have demonstrated the capability to achieve lower FID scores but with higher computation GFLOPS [7]. Recent research highlights the remarkable image and video generation capabilities of diffusion transformer architectures, as demonstrated in methods like Stable Diffusion 3 [8] and Sora [9]. Given the impressive performance of diffusion transformer models, researchers are now training larger and larger DiTs [10]. For instance, Stable Diffusion 3 trained DiT models range from 800M to 8B parameters. Additionally, there is speculation among researchers that Sora might boast around 3 billion parameters. Given the enormous parameter numbers, deploying these DiT models will be costly, especially on certain end devices (e.g., mobile phones).
Figure 2: 2-bit Q-DiT (A) and 2-bit Q-Diffusion (B) quantization results.
@@
To deal with the deployment dilemma, there are recent works on the efficient deployment of diffusion models [11, 12, 13, 14], most of which focus on model quantization. However, as far as we are concerned, there are still two main shortcomings in current research. Firstly, the exploration of quantization methods for transformer-based diffusion models remains quite limited [15, 16] compared to quantizing U-Net-based diffusion models. Secondly, most prevailing quantization approaches heavily rely on post-training quantization (PTQ) [11, 13, 17, 18, 19], which leads to unacceptable performance degradation, particularly with extremely low bit width (e.g., 2-bit and 1-bit). For example, the 2-bit weight quantization results of Q-DiT [15] and Q-Diffusion [11] are shown in Fig. 2. The extremely low-bit quantization of neural networks is essential as it can significantly reduce the computation resources for deployment [20], especially for huge models. During our research, we find that there is still no work considering the extremely low-bit quantization of DiT models.
@@
To tackle these shortcomings, we propose leveraging the quantization-aware training (QAT) technique for the extremely low-bit quantization of large-scale DiT models. Low-bit QAT methods for large-scale models have been discussed in the LLM domain. Recent works have observed that training large language models with extremely low-bit parameters (e.g., binary and ternary) from scratch can also lead to competitive performance [21, 22] compared to their full-precision counterparts. Furthermore, there have been recent advancements in deploying ternary LLMs on CPUs [23], resulting in accelerated inference and reduced energy consumption. These findings highlight the promising potential of ternary quantization and suggest that significant precision redundancy still exists in large-scale models, making QAT a feasible approach for large-scale DiTs.
@@
Inspired by the ongoing in-depth research on ternary LLMs [22, 23], in this paper, we introduce the first attempt at ternary quantization [24] for the DiT model and provide TerDiT, the first quantization scheme for extremely low-bit DiTs to our best knowledge. Our method achieves quantization-aware training (weight only) and efficient deployment for ternary diffusion transformer models. Different from the naive quantization of linear layers in LLMs and CNNs [22, 24], we find that the direct weight ternarization of the adaLN module [25] in DiT blocks [7, 26] leads to large dimension-wise scale and shift values in the normalization layer compared with full-precision models (due to weight quantization, gradient approximation, etc), which results in slower convergence speed and poor model performance. Consequently, we propose a variant of adaLN by applying an RMS Norm [27] after the ternary linear layers of the adaLN module to mitigate this training issue effectively.
@@
With this modification, we scale the ternary DiT model from 600M (size of DiT-XL/2 [7]) to 4.2B (size of Large-DiT-4.2B [28]), with image resolution from 256$\times$256 to 512$\times$512. We further deploy the trained ternary DiT models with 2-bit CUDA kernels, resulting in over 10×10$\times$10 × reduction in checkpoint size and about 8×8$\times$8 × reduction in inference memory consumption while achieving competitive (or even better) generation quality compared with full-precision models. The contributions of our work are summarized as follows:
@@
1) Inspired by the quantization-aware training scheme for low-bit LLMs, we study the QAT method for ternary DiT models and introduce DiT-specific techniques for achieving extremely low-bit quantization for DiT models.
@@
2) We scale ternary DiT models from 600M to 4.2B parameters, with resolution from 256$\times$256 to 512$\times$512, and further deploy the trained ternary DiT on GPUs based on 2-bit CUDA kernels, enabling the inference of a 4.2B DiT with less than 2GB GPU memory (256$\times$256 resolution).
@@
3) Competitive evaluation results compared with full-precision models and baseline quantization methods on the ImageNet [29] benchmark (image generation) showcase the effectiveness of TerDiT.
@@
TerDiT is the first attempt to explore the extremely low-bit quantization of DiT models. We focus on quantization-aware training and efficient deployment for large ternary DiT models, offering valuable insights for future research on deploying extremely low-bit DiT models.
II Related Works
II-A Diffusion Models
@@
Diffusion models have gained significant attention in recent years due to their ability to generate high-quality images and their potential for various applications. The diffusion model was first introduced by [30], proposing a generative model that learns to reverse a diffusion process. This work laid the foundation for subsequent research in the field. [1] further extended the idea by introducing denoising diffusion probabilistic models (DDPMs), which have become a popular choice for image generation tasks. DDPMs have been applied to a wide range of domains, including unconditional image generation [1], image inpainting [31], and image super-resolution [32]. Additionally, diffusion models have been used for text-to-image synthesis, as demonstrated by the DALL-E model [33] and the Imagen model [5]. These models showcase the ability of diffusion models to generate highly realistic and diverse images from textual descriptions. Furthermore, diffusion models have been extended to other modalities, such as audio synthesis [34], video generation [35], and 3D generation [36], demonstrating their versatility and potential for multimodal applications.
@@
II-B Quantization of Diffusion Models
The quantization of diffusion models has been studied in recent years to improve the efficiency of diffusion models. Post-training quantization (PTQ) methods, such as those presented in [11, 13, 17, 18, 19, 15, 16], offer advantages in terms of quantization time and data usage. However, these methods often result in suboptimal performance when applied to low-bit settings. To address this issue, [37] proposes combining quantization-aware low-rank adapters (QALoRA) with PTQ methods, leading to improved evaluation results. As an alternative to PTQ, quantization-aware training (QAT) methods have been introduced specifically for low-bit diffusion model quantization [14, 38, 39, 40]. Despite their effectiveness, these QAT methods are currently only limited to small-sized U-Net-based diffusion models, revealing a research gap in applying QAT to large-scale DiT models. Further exploration of QAT techniques for large DiT models with extremely low bit width could potentially unlock even greater efficiency gains and enable the effective deployment of diffusion models in resource-constrained environments.
II-C Ternary Weight Networks
Ternary weight networks [24] have emerged as a memory-efficient and computation-efficient network structure, offering the potential for significant reductions in inference memory usage. When supported by specialized hardware, ternary weight networks can also deliver substantial computational acceleration. Among quantization methods, ternary weight networks have garnered notable attention, with two primary approaches being explored: weight-only quantization and weight-activation quantization. In weight-only quantization, as discussed in [41], solely the weights are quantized to ternary values. On the other hand, weight-activation quantization, as presented in [42, 43], involves quantizing both the weights and activations to ternary values. Recent research has demonstrated the applicability of ternary weight networks to the training of large language models [22], achieving results comparable to their full-precision counterparts. Moreover, recent developments have enabled the deployment of ternary LLMs on CPUs [23], leading to fast inference and low energy consumption. Building upon these advancements, our work introduces, for the first time, quantization-aware training and efficient deployment schemes specifically designed for ternary DiT models. By leveraging the benefits of ternary quantization in the context of DiT models, we aim to push the boundaries of efficiency and enable the deployment of powerful diffusion models in resource-constrained environments, opening up new possibilities for practical applications.
III TerDiT
In this section, we introduce TerDiT, a framework designed to conduct weight-only quantization-aware training and efficient deployment of large-scale ternary DiT models. We first give a brief review of diffusion transformer (DiT) models in Sec. III-A. Then, building upon the previous open-sourced Large-DiT [28], we illustrate the quantization function and quantization-aware training scheme in Sec. III-B, conduct QAT-specific model structure improvement for better network training in Sec. III-C, and introduce the ternary deployment scheme in Sec. III-D.
III-A Preliminary: Diffusion Transformer Models
Diffusion Transformer. Diffusion transformer [7] (DiT) is an architecture that replaces the commonly used U-Net backbone in the diffusion models with a transformer that operates on latent patches. Similar to the Vision Transformer (ViT) architecture shown in Fig. 3 (C), DiT first patches the spatial inputs into a sequence of tokens, and then the denoising process is carried out through a series of transformer blocks (Fig. 3 (B)). To deal with additional conditional information (e.g., noise timesteps $t$, class labels $l$, natural language inputs), DiT leverages adaptive normalization modules [44] (adaLN-Zero) to insert these extra conditional inputs to the transformer blocks. After the final transformer block, a standard linear decoder is applied to predict the final noise and covariance. The DiT models can be trained in the same way as U-Net-based diffusion models.
@@ III-A Preliminary: Diffusion Transformer Models
AdaLN is an important component in the DiT model [7]. Within the DiT architecture, the adaLN module integrates an MLP layer with a substantial number of parameters, constituting approximately 10% to 20% of the model’s total parameters. Throughout the training of TerDiT, we observe that the direct weight-ternarization of this module yields undesirable training results (analyzed in Sec. III-C).
III-B Model Quantization
As illustrated in Sec. I, there is an increasing popularity in understanding the scaling law of DiT models, which has been proven crucial for developing and optimizing LLMs. In recent explorations, Large-DiT [28] successfully scales up the model parameters from 600M to 7B by incorporating the methodologies of LLaMA [45, 46] and DiT. The results demonstrate that parameter scaling can potentially enhance model performance and improve convergence speed for the label-conditioned ImageNet generation task. Motivated by this, we propose to further investigate the ternarization of DiT models, which can alleviate the challenges associated with deploying large-scale DiT models. In this subsection, we introduce the quantization function and quantization-aware training scheme.
@@ III-B Model Quantization
Referring to current popular quantization methods for LLMs [47, 48], we also multiply a learnable scaling parameter $\alpha$ (randomly initialized) to each ternary linear matrix after quantization, leading to the final value set as $\{-\alpha,0,+\alpha\}$. The quantization function is formulated as:
@@
$\text{RoundClip}(x,a,b)$ $=\text{Clamp}(\text{round}(x),a,b),$ (3)
$\gamma$ $=\frac{1}{mn}\sum_{ij}|W_{ij}|.$ (4)
TerDiT is a weight-only quantization scheme, and we do not quantize the activations.
Quantization-aware Training Scheme. Based on the above-designed quantization function, we train a DiT model from scratch [^1] It is observed in [22] that for ternary LLMs, the conversion or post-training quantization from trained LLMs does not help. So we also train ternary DiT models from scratch. utilizing the straight-through estimator (STE) [49], allowing gradient propagation through the undifferentiable network components. We preserve the full-precision parameters of the network throughout the training process. For each training step, ternary weights are calculated from full-precision parameters by the ternary quantization function in the forward pass, and the gradients of ternary weights are directly applied to the full-precision parameters for parameter updates in the backward pass.
@@
III-C QAT-specific Model Structure Improvement
Activation Analysis for Ternary Linear Layer. In a ternary linear layer, all the parameters take one value from {-$\alpha$, 0, +$\alpha$}. The values passing through this layer would become large activation values, hampering the stable training of neural networks. We conduct a pilot study to qualitatively demonstrate the impact of ternary linear weights on activation values.